OpenStudy (anonymous):

f(x)=5x+4/x+3 and g(x)=3x-5-x find f(g(x)),

OpenStudy (abb0t):

Plug in the whole function into: \(\large \frac{5 \sf \color{red}{x}+4}{\sf \color{red}{x}+3}\)

OpenStudy (abb0t):

those red x's means you plug in the WHOLE function: \(\sf \color{red}{3x-5-x}\) in there.

OpenStudy (anonymous):

place the x by 3x-5-x in f(x) then u should get f(gx)=(5(2x-5)+4)/((2x-5)+4) i.e,(3x-5-x=2x-5) f(gx)=(10x-21)/(2x-1)

OpenStudy (anonymous):

replace not place